PART A: APPLYING THE NORMAL DISTRIBUTION

  1. P( z < -1.42)
  2. P( z > -2.17)
  3. P(z > 2.35)
  4. P(-2.33 < z < 1.84)
  1. 2010 MCAT exam scores are normally distributed with a mean score of 25 with a standard deviation of 6.4
  1. What proportion of students had an MCAT score over 30?
  1. What proportion of students had scores less than 40?
  1. For model year 2010 vehicles the combined city/hwy gas mileage is approximately normal with a mean of 20.3 mpg and a standard deviation of 4.3 mpg.
  1. The 2010 Chevy Camaro gets 19 mpg. What percentage of other 2010 vehicles get better mpg than the Camaro?

  1. How low must a 2010 vehicle’s gas mileage be to fall in the lowest 10% of all vehicles?
  1. What gas mileage would a vehicle have to exceed to be in the top 8% of all 2010 vehicles?
Answer #1

# X - 2010 MCAT Exam Score u=25c 2614 2= 6-73 (a) pla> 30)=P(2>0.78) = 0.2177 The percentane is (2679 (21.77%) (6) Plac40) =